# The geometry of preperiodic points in families of maps on $\mathbb{P}^{N}$

### v1

@@ 2.2. Families of abelian varieties
for $x\in\mathcal{A}_{s}$. The generic Betti rank $\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})$ is the maximal rank of the differential of the Betti map over all points in $\mathcal{X}\cap\pi^{-1}(\Delta)$, and it is independent of the choice of simply connected $\Delta$. Associated to the Betti map and a choice of polarization $\mathcal{L}$ is a $(1,1)$-form $\omega_{\mathcal{A},\mathcal{L}}$ on $\mathcal{A}$, called the Betti form and first introduced by Mok [Mo]; see [DGH2, Proposition 2.2]. It can be defined by
$\omega_{\mathcal{A},\mathcal{L}}=\sum_{1\leq i<j\leq 2g}a_{ij}\,d\beta_{i}\wedge d\beta_{j},$
@@
To see that Conjecture 2.3 is a generalization of Theorem 2.4, let us first replace $S$ with a Zariski-open subset so that all fibers of $\mathcal{A}\to S$ are (smooth) abelian varieties. Consider the multiplication-by-$M$ morphism
$[M]:\mathcal{A}\to\mathcal{A}$
for any choice of integer $M\geq 2$. This endomorphism is polarizable with degree $M^{2}$, and thus induces an embedding
@@
The Green current $\hat{T}_{\Phi_{M,\mathcal{A}}}$ on $S\times\mathbb{P}^{N}$ restricts to the Betti form $\omega_{\mathcal{A},\mathcal{L}}$ on $\mathcal{A}$ associated to the polarization $\mathcal{L}$; indeed, the current is uniquely determined (in its cohomology class $c_{1}(\mathcal{L})$) by its invariance under pullback for the restriction $\Phi_{M,\mathcal{A}}|\mathcal{A}$, where $[M]^{*}\omega_{\mathcal{A},\mathcal{L}}=M^{2}\omega_{\mathcal{A},\mathcal{L}}$. It is then an immediate consequence of [DGH2, Proposition 2.2 (iii)] that
$\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})=2\operatorname{rank}_{\Phi_{M,\mathcal{A}}}(\mathcal{X}).$
It remains to relate the hypothesis on $\mathcal{X}$, that its $\mathbb{Z}$-orbit is Zariski dense in $\mathcal{A}$, to our notion of relative special dimension. Indeed, the condition on $\mathcal{X}\subset\mathcal{A}$ implies that $\mathcal{X}$ is not contained in any proper subgroup scheme over $S$ nor a torsion-translate of such a subgroup. In particular, the smallest $\Phi_{M,\mathcal{A}}$-special subvariety containing $\mathcal{X}$ in $S\times\mathbb{P}^{N}$ is the embedded copy of $\mathcal{A}$ itself. That is, $r_{\Phi,\mathcal{X}}=g$, the relative dimension of $\mathcal{A}$.
@@
Comparing the statement of Theorem 2.4 with Conjectures 1.1 and 2.3, it is important to notice some additional complications that arise in the dynamcial setting. It can happen that the $\Phi$-orbit of a subvariety $\mathcal{X}\subset S\times\mathbb{P}^{N}$ is Zariski dense in some $\Phi$-special subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$, while $r_{\Phi,\mathcal{X}}<\dim_{S}\mathcal{Y}$. Such examples arose in the formulation of the Dynamical Manin-Mumford conjecture; see [GTZ] for details. Those examples came from abelian varieties, where the $\Phi$ in question is built from complex multiplication, while the smaller $\Phi$-special is invariant for the usual $\mathbb{Z}$-action.
@@ 2.3. Non-degeneracy
We conclude this section by observing that there is an important notion in the literature which is similar to but distinct from condition (2) of Conjecture 1.1; we include it here for comparison. By definition, $\mathcal{X}$ in $S\times\mathbb{P}^{N}$ is non-degenerate for $\Phi$ if $\hat{T}_{\Phi}^{\wedge\dim\mathcal{X}}\wedge[\mathcal{X}]\not=0$. This concept was introduced for subvarieties $\mathcal{X}$ in families of powers of elliptic curves by Habegger [Ha] and played an important role in the proof of the Geometric Bogomolov Conjecture over characteristic 0 by Gao–Habegger and Cantat–Gao–Habegger–Xie [GH1, CGHX], the proof of the uniform Mordell–Lang conjecture by Gao–Habegger, Kühne and Gao–Ge–Kühne [DGH1, DGH2, , GGK] and the proof of the relative Manin-Mumford conjecture by Gao–Habegger [GH2]. In [YZ1], this notion of non-degeneracy was extended to families of polarized dynamical systems, and it is a key hypothesis for their theorems on arithmetic equidistribution. In particular, if $\Phi$ and $\mathcal{X}$ are defined over $\overline{\mathbb{Q}}$, then Conjecture 1.1 would characterize the existence of a Zariski-dense set of small points in $\mathcal{X}$, while the non-degeneracy condition would imply that the Galois orbits of these small points are uniformly distributed with respect to the measure $\hat{T}_{\Phi}^{\wedge\dim\mathcal{X}}\wedge[\mathcal{X}]\not=0$ (if the points exist) [YZ1, Theorem 5.4.3] [Ga2]. Note that every projective algebraic subvariety of $\mathbb{P}^{N}$ is non-degenerate for a morphism $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$ when $\dim S=0$, by Lemma 2.3.
3. Some known cases of Conjecture 1.1
In this section, we provide examples where Conjecture 1.1 is known, moving away from the setting of abelian varieties. We begin with the original Dynamical Manin-Mumford Conjecture, corresponding to the case of $\dim S=0$. In §3.2, we observe that the Conjecture holds in all cases when $N=1$, as it can be deduced from [De3, Theorem 1.1], and we relate it to the theory of $J$-stability for maps on $\mathbb{P}^{1}$. In §3.3 we explain that Conjecture 1.1 implies the so-called Dynamical André-Oort conjecture (or “DAO”). Finally, we illustrate that uniform versions of the conjecture are consequences of the conjecture itself, with the example of shared preperiodic points for distinct maps on $\mathbb{P}^{1}$.
@@ 3.1. The Dynamical Manin-Mumford Conjecture
When $S$ is a point, so that $\dim S=0$ and $\Phi:\mathbb{P}^{N}\to\mathbb{P}^{N}$ is an endomorphism over $\mathbb{C}$, Conjectures 1.1 and 1.2 reduce to Zhang’s Dynamical Manin-Mumford Conjecture (or DMM), as reformulated by Ghioca and Tucker in [GT]: a subvariety $X$ of $\mathbb{P}^{N}$ contains a Zariski-dense set of preperiodic points if and only if it is a $\Phi$-special subvariety. Indeed, the current $T_{\Phi}^{\wedge r}\wedge[X]$ will vanish if $r>\dim X$, but is nonzero for $r=\dim X$ by Lemma 2.1. The implication that a $\Phi$-special subvariety of $\mathbb{P}^{N}$ always contains a dense set of preperiodic points is well known; see for example [Fa1, BD, DS1]. The converse implication has been proved in few settings outside of the cases of endomorphisms of abelian varieties.
It is worth observing that DMM is obvious for maps on $\mathbb{P}^{1}$: a single point is either preperiodic (which is equivalent to special for points) or it is not. The DMM conjecture has been fully resolved for polarized endomorphisms of $(\mathbb{P}^{1})^{N}$ over $\mathbb{C}$ [GNY, GNY, MSW], for subvarieties $X\subset(\mathbb{P}^{1})^{N}$ of arbitrary dimension. Very recently, Dujardin, Favre, and Ruggiero completed a proof for polynomial maps of $\mathbb{A}^{2}$ that extend regularly to $\mathbb{P}^{2}$, where $X$ is a complex algebraic curve in $\mathbb{P}^{2}$ (though there remains a condition on the curve, regarding its intersection with the line at infinity) [DFR].
A related problem for polynomial automorphisms of $\mathbb{A}^{2}$ was treated in [DF]. We do not consider their setting in this article.
@@ 3.2. Dimension $N=1$
Conjecture 1.1 also implies [De3, Theorem 1.1] (and therefore also [DF, Theorem 2.5] and [Mc, Lemma 2.1] addressing the case of marked critical points in $\mathbb{P}^{1}$), so it is logically equivalent in dimension $N=1$. Indeed, suppose that $\Phi:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ is an algebraic family of maps of degree $d>1$, and suppose that $\Gamma_{a}\subset S\times\mathbb{P}^{1}$ is the graph of a marked point $a:S\to\mathbb{P}^{1}$. Assume the pair $(\Phi,a)$ is stable, meaning that $\hat{T}_{\Phi}\wedge[\Gamma_{a}]=0$. We will deduce from Conjecture 1.1 that $a$ is either persistently preperiodic or the pair $(\Phi,a)$ is isotrivial. It suffices to assume that $\dim S=1$. Passing to a branched cover of $S$ if necessary, we can mark three distinct periodic points for $\Phi$ and, removing a Zariski-closed subset of $S$ where they collide, we can change coordinates on $\mathbb{P}^{1}$ so that $\{0,1,\infty\}$ are persistently periodic. Stability of $(\Phi,a)$ implies one of two things: either (1) the graph $\Gamma_{a}$ is itself preperiodic (i.e., the special dimension is $r_{\Phi,\Gamma_{a}}=0$), or (2) as a consequence of Conjecture 1.1 with $r_{\Phi,\Gamma_{a}}=1$, the preperiodic points of $\Phi$ cannot be Zariski dense in $\Gamma_{a}$. In case (1), we are done. In case (2), this means that there is a Zariski-open subset $U$ of $S$ over which the point $a$ is never preperiodic. In particular, the point $a$ and its infinite forward orbit is disjoint from the set $\{0,1,\infty\}$ over the quasiprojective curve $U$. But, as McMullen observed in his proof of [Mc, Lemma 2.1], there are at most finitely many non-constant holomorphic functions $U\to\mathbb{P}^{1}\setminus\{0,1,\infty\}$. This implies that the iterates $\Phi^{n}(a)$ must be constant (and distinct) over $U$, for all $n$ sufficiently large. But then, by a simple interpolation argument, we see that $\Phi_{s}$ is independent of the parameter $s\in U$. In other words, in case (2), the family $\Phi$ is isotrivial.
@@
is the critical locus. (This $\mathcal{X}$ is not necessarily irreducible, but we can apply Theorem 3.1 to each component.) Then $r_{\Phi,\mathcal{X}}=0$ (for all components) if and only if $\Phi$ is a family of postcritically finite maps. We know from Thurston rigidity that this can hold if and only if $\Phi$ is either an isotrivial family or a family of flexible Lattès maps [DH].
@@
projects to the parameter space $S$ as the bifurcation current $T_{\mathrm{bif}}$ for the family $\Phi$. See [De1] and [DF] for definitions. It was proved in [De1] that the bifurcation current vanishes if and only if the family is $J$-stable in the sense of Mañé-Sad-Sullivan [MSS] and Lyubich [Ly]. So Conjecture 1.1 (i.e., Theorem 3.1) implies the (known) characterization of instability in algebraic families $\Phi$ by the existence of many non-persistent preperiodic critical points; compare [Mc, Theorem 2.2] and [DF, Theorem 2.5].
@@ 3.3. DAO as a special case
Let $f:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ be an algebraic family of rational maps on $\mathbb{P}^{1}$ of degree $d\geq 2$. We say it has dimension $m$ in moduli if the induced projection from $S$ to $\mathrm{M}_{d}^{1}$, the moduli space of all maps of degree $d$ on $\mathbb{P}^{1}$, has $m$-dimensional image. Here we observe that Conjecture 1.1 (in fact, in its weaker form of Conjecture 1.2) implies the following conjecture proposed by Baker-DeMarco [BD2] and Ghioca-Hsia-Tucker [GHT1]; the statement here is given explicitly in [De3, Conjecture 6.1].
@@
By definition, marked points $(a_{0},\ldots,a_{m})$ are dynamically related along $S$ if there exists a proper subvariety $\mathcal{Y}\subset S\times(\mathbb{P}^{1})^{m+1}$ projecting dominantly to $S$ which is preperiodic for the fiber-product $\Phi=f^{[m+1]}$ defined by
@@
To place Conjecture 3.4 in the setting of Conjectures 1.1 and 1.2, note that the $\Phi$ of (3.1) defines a family of polarized endomorphisms of $(\mathbb{P}^{1})^{m+1}$ and so extends to a family of endomorphisms of projective space $\mathbb{P}^{M}$ for some dimension $M\geq m+1$. In particular, the collection $(a_{0},\ldots,a_{m})$ is dynamically related if and only if the graph $\Gamma$ of this $(m+1)$-tuple lies in a $\Phi$-special subvariety of relative dimension $\leq m$ over $S$. Indeed, for powers of nonisotrivial maps on $\mathbb{P}^{1}$ such as $\Phi$, all $\Phi$-special subvarieties in $S\times(\mathbb{P}^{1})^{m+1}$ will be preperiodic for $\Phi$; see, for example, the discussion in Section 2 of [GT], as Lattès maps coming from complex multiplication are rigid in $\mathrm{M}_{d}^{1}$.

### v2

@@ 2.2. Families of abelian varieties
for $x\in\mathcal{A}_{s}$. The generic Betti rank $\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})$ is the maximal rank of the differential of the Betti map over all smooth points in $\mathcal{X}\cap\pi^{-1}(\Delta)$, and it is independent of the choice of simply connected $\Delta$. Associated to the Betti map and a choice of polarization $\mathcal{L}$ is a $(1,1)$-form $\omega_{\mathcal{A},\mathcal{L}}$ on $\mathcal{A}$, called the Betti form and first introduced by Mok [Mo]; see [DGH2, Proposition 2.2]. It can be defined by
(2.1) $\omega_{\mathcal{A},\mathcal{L}}=\sum_{1\leq i<j\leq 2g}a_{ij}\,d\beta_{i}\wedge d\beta_{j},$
@@
Proposition 2.5.
Conjecture 2.3 implies Theorem 2.4.
Proof.
Let us first replace $S$ with a Zariski-open subset so that all fibers of $\mathcal{A}\to S$ are (smooth) abelian varieties. Consider the multiplication-by-$M$ morphism
$M]:\mathcal{A}\to\mathcal{A}$
for any choice of integer $M\geq 2$. The endomorphism is polarizable with degree $M^{2}$ on each fiber of the projection $\mathcal{A}\to S$. Again replacing $S$ with a Zariski-open subset if necessary, there is an embedding
@@
The Green current $\hat{T}_{\Phi_{M,\mathcal{A}}}$ on $S\times\mathbb{P}^{N}$ restricts to the Betti form $\omega_{\mathcal{A},\mathcal{L}}$ on $\mathcal{A}$ associated to the polarization $\mathcal{L}$; indeed, the current is uniquely determined (in its cohomology class $c_{1}(\mathcal{L})$) by its invariance under pullback for the restriction $\Phi_{M,\mathcal{A}}|_{\mathcal{A}}$, where $M]^{*}\omega_{\mathcal{A},\mathcal{L}}=M^{2}\omega_{\mathcal{A},\mathcal{L}}$.
As in the proof of [DGH2, Proposition 2.2 (iii)], we have
(2.2) $\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})=2\operatorname{rank}_{\Phi_{M,\mathcal{A}}}(\mathcal{X}).$
In detail, let $2k$ be the generic Betti rank of $\mathcal{X}$; it is always even by, for example, the formulas provided in [ACZ]. It is clear that $\operatorname{rank}_{\Phi_{M,\mathcal{A}}}(\mathcal{X})\leq k$ from the definition of the Betti form $\omega_{\mathcal{A},\mathcal{L}}$ given in (2.1). On the other hand, as pointed out in [CGHX, §2.4], the Betti form acts on pairs of tangent vectors in $\mathcal{A}$ by the composition of a complex-linear projection to the tangent space to a fiber $\mathcal{A}_{s}$ and then applying the Kähler form $\omega_{s}=\omega_{\mathcal{A},\mathcal{S}}|_{\mathcal{A}_{s}}$. In particular, since the image under the projection from $T_{x}\mathcal{X}$ to $T_{x}\mathcal{A}_{s}$ at a point $x\in\mathcal{X}$ is generally a complex subspace of dimension $k$, we deduce that $\omega_{\mathcal{A},\mathcal{L}}^{\wedge k}\not=0$. In other words, $\operatorname{rank}_{\Phi_{M,\mathcal{A}}}(\mathcal{X})\geq k$, demonstrating equality in (2.2).
It remains to relate the hypothesis on $\mathcal{X}$, that its $\mathbb{Z}$-orbit is Zariski dense in $\mathcal{A}$, to our notion of relative special dimension. Indeed, the condition on $\mathcal{X}\subset\mathcal{A}$ implies that $\mathcal{X}$ is not contained in any proper subgroup scheme over $S$ nor a torsion-translate of such a subgroup. In particular, the smallest $\Phi_{M,\mathcal{A}}$-special subvariety containing $\mathcal{X}$ in $S\times\mathbb{P}^{N}$ is the embedded copy of $\mathcal{A}$ itself. That is, $r_{\Phi,\mathcal{X}}=g$, the relative dimension of $\mathcal{A}$. ∎
@@
Comparing the statement of Theorem 2.4 with Conjectures 1.1 and 2.3, it is important to note some additional complications that arise in the dynamical setting. It can happen that the $\Phi$-orbit of a subvariety $\mathcal{X}\subset S\times\mathbb{P}^{N}$ is Zariski dense in some $\Phi$-special subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$, while $r_{\Phi,\mathcal{X}}<\dim_{S}\mathcal{Y}$. Such examples led to the formulations of the Dynamical Manin-Mumford conjecture in [GTZ] and [GT] and the introduction of the auxiliary endomorphism $\Psi$ in the definition of $\Phi$-special. For example, $\Phi:E\times E\to E\times E$ can act by a product of complex-multiplication endomorphisms on an elliptic curve $E$, chosen so that the diagonal $\Delta_{E}\subset E\times E$ (which contains a Zariski-dense set of preperiodic points, being the torsion points of $E$) is not a preperiodic curve for $\Phi$, so its $\Phi$-orbit is Zariski-dense in $E\times E$. On the other hand, the diagonal $\Delta_{E}$ invariant under the usual $\mathbb{Z}$-action on $E\times E$, so $r_{\Phi,\Delta_{E}}=1$. See [GTZ] and [Paz] for details and explicit constructions.
@@ 2.3. Non-degeneracy
We conclude this section by observing that there is an important notion in the literature which is similar to but distinct from condition (2) of Conjecture 1.1; we include it here for comparison. By definition, $\mathcal{X}$ in $S\times\mathbb{P}^{N}$ is non-degenerate for $\Phi$ if $\hat{T}_{\Phi}^{\wedge\dim\mathcal{X}}\wedge[\mathcal{X}]\not=0$. This concept was introduced for subvarieties $\mathcal{X}$ in families of powers of elliptic curves by Habegger [Ha] and played an important role in the proof of the Geometric Bogomolov Conjecture over function fields in characteristic 0 by Gao–Habegger and Cantat–Gao–Habegger–Xie [GH1, CGHX], the proof of the uniform Mordell–Lang conjecture by Gao–Habegger, Kühne and Gao–Ge–Kühne [DGH1, DGH2, , GGK] and the proof of the relative Manin-Mumford conjecture by Gao–Habegger [GH2]. In [YZ1], this notion of non-degeneracy was extended to families of polarized dynamical systems, and it is a key hypothesis for their theorems on arithmetic equidistribution. In particular, if $\Phi$ and $\mathcal{X}$ are defined over a number field $K$, then Conjecture 1.1 would characterize the existence of a Zariski-dense set of small points in $\mathcal{X}$, while the non-degeneracy condition would imply that the $\operatorname{Gal}(\overline{K}/K)$-orbits of these small points are uniformly distributed with respect to the measure $\hat{T}_{\Phi}^{\wedge\dim\mathcal{X}}\wedge[\mathcal{X}]\not=0$ (if the points exist), as proved in [YZ1, Theorem 6.2.3] [Ga2, Theorem 2] [GTV, Theorem 6.2]. Note that every projective algebraic subvariety of $\mathbb{P}^{N}$ is non-degenerate for a morphism $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$ when $\dim S=0$, by Lemma 2.3.
3. Some special cases of Conjecture 1.1
In this section, we provide examples where Conjecture 1.1 is known, moving away from the setting of abelian varieties. We begin in the setting of the original Dynamical Manin-Mumford Conjecture, corresponding to the case of $\dim S=0$, in §3.1. In §3.2, we observe that Conjecture 1.1 holds in all cases when $N=1$, and we relate it to the theory of $J$-stability for maps on $\mathbb{P}^{1}$. In §3.3 we explain that Conjecture 1.1 implies the so-called Dynamical André-Oort conjecture (or “DAO”). Finally, in §3.4 we illustrate that uniform versions of the conjecture are consequences of the conjecture itself, with the example of shared preperiodic points for distinct maps on $\mathbb{P}^{1}$.
@@ 3.1. The Dynamical Manin-Mumford Conjecture
When $S$ is a point, so that $\dim S=0$ and $\Phi:\mathbb{P}^{N}\to\mathbb{P}^{N}$ is an endomorphism over $\mathbb{C}$, Conjectures 1.1 and 1.3 reduce to Zhang’s Dynamical Manin-Mumford Conjecture (or DMM), as reformulated by Ghioca and Tucker in [GT]: conjecturally, a subvariety $X$ of $\mathbb{P}^{N}$ contains a Zariski-dense set of preperiodic points if and only if it is a $\Phi$-special subvariety. The reduction follows from observing that the current $T_{\Phi}^{\wedge r}\wedge[X]$ will vanish if $r>\dim X$, but is nonzero for $r=\dim X$ by Lemma 2.1. The implication that a $\Phi$-special subvariety of $\mathbb{P}^{N}$ always contains a dense set of preperiodic points is well known; see for example [Fa1, BD, DS1]. However, the converse implication has been proved only in a few settings outside of the cases of endomorphisms of abelian varieties, all of which we outline here.
It is worth observing that DMM is obvious for maps on $\mathbb{P}^{1}$: a single point is either preperiodic (which is equivalent to special for points) or it is not. For $N>1$, the DMM conjecture has been fully resolved for polarized endomorphisms of $(\mathbb{P}^{1})^{N}$ over $\mathbb{C}$ [GNY2, GNY1, MSW], for subvarieties $X\subset(\mathbb{P}^{1})^{N}$ of arbitrary dimension. A proof of DMM for polynomial maps of $\mathbb{A}^{2}$ that extend regularly to $\mathbb{P}^{2}$, assuming that the complex algebraic curve $X$ satisfies a certain condition on its intersection with the line at infinity, was provided in [DFR]. A related problem for polynomial automorphisms of $\mathbb{A}^{2}$ was treated in [DF], but Conjecture 1.1 does not cover this setting.
@@ 3.2. Dimension $N=1$
Conjecture 1.1 also implies [De3, Theorem 1.1] (and therefore also [DF, Theorem 2.5] and [Mc, Lemma 2.1] addressing the case of marked critical points in $\mathbb{P}^{1}$), so it is logically equivalent in dimension $N=1$. Indeed, suppose that $\Phi:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ is an algebraic family of maps of degree $d>1$, and suppose that $\Gamma_{a}\subset S\times\mathbb{P}^{1}$ is the graph of a marked point $a:S\to\mathbb{P}^{1}$. Assume the pair $(\Phi,a)$ is stable, so that $\hat{T}_{\Phi}\wedge[\Gamma_{a}]=0$. We will deduce from Conjecture 1.1 that $a$ is either persistently preperiodic or the pair $(\Phi,a)$ is isotrivial. It suffices to assume that $\dim S=1$. Passing to a branched cover of $S$ if necessary, we can mark three distinct periodic points for $\Phi$ and, removing a Zariski-closed subset of $S$ where they collide, we can change coordinates on $\mathbb{P}^{1}$ so that $\{0,1,\infty\}$ are persistently periodic. Stability of $(\Phi,a)$ implies one of two things: either (1) the graph $\Gamma_{a}$ is itself preperiodic (i.e., the special dimension is $r_{\Phi,\Gamma_{a}}=0$), or (2) as a consequence of Conjecture 1.1 with $r_{\Phi,\Gamma_{a}}=1$, the preperiodic points of $\Phi$ cannot be Zariski dense in $\Gamma_{a}$. In case (1), we are done. In case (2), this means that there is a Zariski-open subset $U$ of $S$ over which the point $a$ is never preperiodic. In particular, the point $a$ and its infinite forward orbit is disjoint from the set $\{0,1,\infty\}$ over the quasiprojective curve $U$. But, as McMullen observed in his proof of [Mc, Lemma 2.1], there are at most finitely many non-constant holomorphic functions $U\to\mathbb{P}^{1}\setminus\{0,1,\infty\}$. This implies that the iterates $\Phi^{n}(a)$ must be constant (and distinct) over $U$, for all $n$ sufficiently large. But then, by a simple interpolation argument, we see that $\Phi_{s}$ is independent of the parameter $s\in U$. In other words, in case (2), the family $\Phi$ is isotrivial.
@@
is the critical locus. (This $\mathcal{X}$ is not necessarily irreducible, but we can apply Theorem 3.1 to each component.) Then $r_{\Phi,\mathcal{X}}=0$ (for all components) if and only if $\Phi$ is a family of postcritically finite maps. We know from Thurston rigidity that this can hold if and only if $\Phi$ is either an isotrivial family or a family of flexible Lattès maps; see [DH, Theorem 1] or [Mc, Theorem 6.2].
@@
projects to the parameter space $S$ as the bifurcation current $T_{\mathrm{bif}}$ for the family $\Phi$. See [De1] and [DF] for definitions. It was proved in [De1] that the bifurcation current vanishes if and only if the family is $J$-stable in the sense of Mañé-Sad-Sullivan [MSS] and Lyubich [Ly]. So Conjecture 1.1 (i.e., Theorem 3.1) implies a (known) characterization of instability in algebraic families $\Phi$ by the existence of many parameters with (non-persistent) preperiodic critical points; compare [Mc, Lemma 2.1] and [DF, Theorem 2.5].
@@ 3.3. DAO as a special case
Let $f:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ be an algebraic family of rational maps on $\mathbb{P}^{1}$ of degree $d\geq 2$. We say the family has dimension $m$ in moduli if the induced projection from $S$ to $\mathrm{M}_{d}^{1}$, the moduli space of all maps of degree $d$ on $\mathbb{P}^{1}$ modulo Möbius conjugacy, has $m$-dimensional image. Here we observe that Conjecture 1.1 (in fact, in its weaker form of Conjecture 1.3) implies the following conjecture proposed by Baker-DeMarco [BD2] and Ghioca-Hsia-Tucker [GHT1]; the statement here is given explicitly in [De3, Conjecture 6.1].
@@
By definition, marked points $(a_{0},\ldots,a_{m})$ are dynamically related along $S$ if there exists a proper closed subvariety $\mathcal{Y}\subset S\times(\mathbb{P}^{1})^{m+1}$ projecting dominantly to $S$ which is preperiodic for the fiber-product $\Phi=f^{[m+1]}$ defined by
@@
To place Conjecture 3.4 in the setting of Conjectures 1.1 and 1.3, note that the $\Phi$ of (3.1) defines a family of polarized endomorphisms of $(\mathbb{P}^{1})^{m+1}$ and so extends to a family of endomorphisms of projective space $\mathbb{P}^{M}$ for some dimension $M\geq m+1$. In particular, the collection $(a_{0},\ldots,a_{m})$ is dynamically related if and only if the graph $\Gamma$ of this $(m+1)$-tuple lies in a $\Phi$-special subvariety of relative dimension $\leq m$ over $S$. Indeed, for powers of nonisotrivial maps on $\mathbb{P}^{1}$ such as $\Phi$, all $\Phi$-special subvarieties in $S\times(\mathbb{P}^{1})^{m+1}$ will be preperiodic for $\Phi$; see, for example, the discussion in Section 2 of [GT], as Lattès maps coming from complex multiplication are rigid in $\mathrm{M}_{d}^{1}$.